OpenStudy (anonymous):

Please Help! Giving Medal A plumber needs a certain part to complete a repair job. The probability that one part supplier has the part is 0.81, the probability that the other part supplier has the part is 0.90, and the probability that both have the part is 0.79. Find the probability that the plumber will be able to complete the repair.

OpenStudy (anonymous):

Pr(A or B) = Pr(A) + Pr(B) - Pr(A and B) P(supplier 1 or supplier 2 has it) = Pr(1 has it) + Pr(2 has it) - Pr (both have it) 0.81 + 0.90 - 0.79 = 0.92

OpenStudy (anonymous):

Thank You =)